A spring extends 3 inches when loaded vertically with 4 lbs. This weight was pulled further down by 3 inches an was released from rest. Determine the frequency of vibration in Hz


1
Expert's answer
2021-04-21T07:04:28-0400

You

15:08


"extension of spring = 3 inches = 0.0762 Meters"

"load = 4lbs = 1.814kgs" =18.14N

"k, spring constant = load\/ext"

"=18.4\/0.0762"

"=" 238.05N/m

weight required to pull the spring a further 3inches (0.0762m)

"=238.05*2(0.0762)"

"=36.27882N"

frequency F, = "1/2\pi *\sqrt{K/m}"

"=1/2\pi * \sqrt{238.05/36.27882}"

=4.02 Hz.